OpenStudy (wcrmelissa2001):

Help please!! I have a question regarding probability. E.g. the question is like if Jack comes to crossroads, he can either go path A or path B. Probability he goes path A is 1/3. Jack comes to another crossroad. He can go either C or D. PRobabilty of picking C is 1/2 (values aren't right but just go along with it) So what is the probability of him going along path a to c? Do i just take 1/2+1/3?

OpenStudy (sunnnystrong):

He has a probability of going down route b of 66% and route a of 33%. He has a probability of going down route c or d 50%.

OpenStudy (sunnnystrong):

He is more likely to take route b to either c or d. I want to guess and say 35% probability of take b-c and b-d And 15% probability he will take a-c or a-d

OpenStudy (mww):

Use a tree diagram for multi step probability. 

OpenStudy (stacey):

You multiply as @mww shows in the tree diagram. If done correctly, all the probabilities of the possible outcomes will add up to 1. \[\frac{ 1 }{ 6 }+\frac{ 1 }{ 6 }+\frac{ 2 }{ 3 }=1\]

OpenStudy (wcrmelissa2001):

wait so probability of him going from a to c would be like er..1/6?

OpenStudy (killuaxzoldyck):

https://akk.li/pics/anne/jpg

OpenStudy (wcrmelissa2001):

@mww

OpenStudy (mww):

yep multi step is always multiply ALONG the branch. Different outcomes, say the probability of going A to C OR going A to D would be adding the probabilities BETWEEN branches

OpenStudy (wcrmelissa2001):

THANKS :D